OpenStudy (anonymous):

I give medals! What is the perimeter of a rectangle with a width of 20.3 cm and length of 30.2 cm?

OpenStudy (anonymous):

the formula for finding the perimeter of a rectangle is P=2L+2W so P=2(30.2) + 2(20.3) P=60.4 +40.6 P=101